A friend asked for my help, now I need yours
5X(X-1)(2X-9)(8X-3)=0

Answer by venugopalramana(3286)

5X(X-1)(2X-9)(8X-3)=0...a product is zero if any one of the factors is zero.so the eqn.implies
5X=0..or..(X-1)=0..or..(2X-9)=0..or..(8X-3)=0..so we get
x=0...or....x=1.....or...2x=9....or...8x=3.....
x=0...or....x=1.....or...x=9/2....or...x=3/8.....IS THE ANSWER